en block 30cm on a side has a density of0.7o/em. If it is floating in freshwater (which ensity of lg/cm), what percentage of the wood block is above the water level? has a d
Answer #1

Let the height of the block inside the water level be 'h'

Then according to the law of floatation of bodies:

Weight of the body = Weight of the displaced volume

ho_s imes V_s imes g = ho_w imes V_d imes g

0.7 x (30)3 × q 1 × h × 302 × q

1 × h × 302 0.7 × (30)

h =0.7× (30) = 21cm2

Hence the height of body above liquid level = 30 - 21 = 9 cm

So, the percentage of wood block above water level = Vol. of Solid above watrvl 302 x 9 Total Vol 100% 30% 303